I need help to solve this problem 8/x+2+8/x-2=3
Answer by algebrahouse.com(1659)

"I need help to solve this problem 8/x+2 + 8/x-2 = 3"

(x + 2)(x - 2)[8/(x + 2) + 8/(x - 2) = 3] {multiplied entire equation by LCD, (x + 2)(x - 2)}
(x + 2)(x - 2)(8/x + 2) + (x + 2)(x - 2)(8/x - 2) = (x + 2)(x - 2)(3) {multiplied each term by (x + 2)(x - 2)}
8(x - 2) + 8(x + 2) = 3(x^2 - 4) {cancelled and multiplied what's left}
8x - 16 + 8x + 16 = 3x^2 - 12 {used distributive property}
16x = 3x^2 - 12 {combined like terms}
3x^2 - 16x - 12 {subtracted 16x from both sides}
(3x + 2)(x - 6) = 0 {factored into two binomials}
3x + 2 = 0 or x - 6 = 0 {set each factor equal to 0}
3x = -2 or x = 6 {subtracted 2 from both sides in 1st equation and added 6 in 2nd}
x = -2/3 or x = 6 {divided both sides by 3 in 1st equation}
www.algebrahouse.com